OpenStudy (anonymous):

how to use L'Hopital's rule on this problem? lim x -> ∞ (x + e^x)^1/x

OpenStudy (jkristia):

I just learned L'Hopital myself, so I might not be able to help, but do you know how to start ?

OpenStudy (anonymous):

not at all =/

OpenStudy (anonymous):

Hi \[\lim_{x \rightarrow \infty}(x+e^{x})^{1/x} \rightarrow \lim_{x \rightarrow \infty}1/x*\ln(x+e^x)\] \[hopital-->\lim (1+e^x)/(x+e^x)-->hopital-->e^x/e^x=1 \] \[answer=e^1=e\]

OpenStudy (anonymous):

thanks alot but i dont get much

OpenStudy (jkristia):

You need to see if the limit is of the inderterminant form \[\frac{ 0 }{ 0 }, or \frac{ \pm \infty }{ \pm \infty }\] If it is, then you can apply L'Hopital

OpenStudy (jkristia):

or \[\infty^0\] which is the case here

OpenStudy (anonymous):

ok

OpenStudy (jkristia):

so, to bring down the exponent you have to take the ln of both sides (to use the property of logarithm)

OpenStudy (anonymous):

oh ths

OpenStudy (anonymous):

when you use ln the expression becames \[\infty/\infty\] in the last step \[e^{\ln x}=x\]

OpenStudy (anonymous):

ok

OpenStudy (jkristia):

\[\ln y = \lim_{x \rightarrow \infty} (\frac{ 1 }{ x } \ln (1+e^x))\] You still have the form \[\frac{ \infty }{ \infty }\] so now you can apply L'Hopital

OpenStudy (jkristia):

ok - you got your answer

OpenStudy (anonymous):

ok ths

OpenStudy (jkristia):

do you get how to solve it ?

OpenStudy (anonymous):

i thnik so

OpenStudy (anonymous):

i need to write it down step by sstep maybe i will get it all

OpenStudy (anonymous):

wut do i do when i bring down the exponent again

OpenStudy (anonymous):

i got 1/x * ln(x+e^x)

OpenStudy (jkristia):

what is the next step after you get \[\ln (y) = \lim_{x \rightarrow \infty}\left[ \frac{ \ln(1+e^x) }{ x } \right]\]

OpenStudy (jkristia):

right, now you brought down the exponent, o now you have the inderterminant form infinity/infinity So now you can apply L'Hopital, meaning take the derivative of the numerator and the derivative of the denominator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got ln(y) = e^x / x + e^x

OpenStudy (jkristia):

no, check your denominator again

OpenStudy (anonymous):

i got x = 1

OpenStudy (jkristia):

yes, so you get (1 + e^x)(1) = 1 + e^x

OpenStudy (jkristia):

so now you have (e^x)/(1+e^x), that is the inderterminant form of infinity / infinity, so you can apply L'Hopital again

OpenStudy (anonymous):

i took the derivative and got e^x / x + e^x because the derivate of ln is 1/x ?

OpenStudy (jkristia):

let me double check my answer, I think I got it right

OpenStudy (anonymous):

ok

OpenStudy (jkristia):

\[\frac{ d }{ dx } \ln(1 + e^x) = \frac{ \frac{ d }{ dx } (1 + e^x) }{ (1 + e^x) } = \frac{ e^x }{ 1 + e^x }\] Isn't that correct ?

OpenStudy (anonymous):

yes

OpenStudy (jkristia):

and d/dx if x = 1, so you have the denominator (1+e^x)(1), where (1) is the derivative of x

OpenStudy (jkristia):

so \[\ln (y) = \lim_{x \rightarrow \infty}\left[ \frac{ e^x }{ (1+e^x)(1) } \right] = \frac{ \infty }{ \infty }\]

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

I think the problem was x+e^x so you have to write \[d/dx \ln(x+e^x)=(1+e^x)/(x+e^x)\]

OpenStudy (jkristia):

and if you apply L'Hopital again, then you get e^x / e^x which cancels and become 1 Then ln(y) = 1

OpenStudy (jkristia):

ohhhh - sorry - I missed that, I was doing 1 + e^x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks for walking me through, i think i got it

OpenStudy (jkristia):

see, I use (1+..) every where , ok - my mistake, I will go back and redo it :) But the steps should still be the same, if you get the inderterminant form infinity/infinity then apply L'hopital again, at the end you get ln(y) = 'something' then to find y you raise both sides by e

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ths